OpenStudy (anonymous):

sigma (k=1) to infinity (-1)^(k+1) sqrt(k)/k+1 . I tried to solve it using the ratio test, but got stuck in the end, and I can't use root test since the whole thing is not to the power of 1/k. Which way can I use?

OpenStudy (anonymous):

\[\frac{(k+1)\sqrt(k+1)}{k^{3/2} + 2 \sqrt(k)}\] that's where I got stuck after using the ratio test

OpenStudy (anonymous):

\[\sum_{k=1}^{\infty}(-1)^{k+1} \frac{\sqrt(k)}{k+1}\] that's the question

OpenStudy (anonymous):

oh wow, alas, visitors lol ^_^

OpenStudy (anonymous):

It doesn't say what you are doing? test for convergence?

OpenStudy (anonymous):

lol, yes ^_^ but I'm not sure which way to choose, I tried ratio test

OpenStudy (anonymous):

umm it looks like you could maybe use divergence test

OpenStudy (anonymous):

kth term divergence test? ^_^

OpenStudy (anonymous):

so from where I stopped , I can use the k-th term divergence test? right?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

kk ^_^

OpenStudy (anonymous):

ok with divergence test you get = 0 so that is inconclusive

OpenStudy (anonymous):

so i'll try a ratio

OpenStudy (anonymous):

no, so that means it's convergent!

OpenStudy (anonymous):

but, the book says it's conditionally convergent . That's where I'm confused :(

OpenStudy (anonymous):

oh I see now! The series is increasing, but it's limit tends to zero, so that's why it's conditonally convergent right? ^_^

OpenStudy (anonymous):

okay so when you get 0 from the divergence test, that tells you to try another test basically. And if you can get it to converge with another test (that is not absolute convergence) then its conditionally convergent

OpenStudy (anonymous):

um, that's a good guess. But I don't think so

OpenStudy (anonymous):

so with the ratio test I got 1

OpenStudy (anonymous):

so again, that means nothing

OpenStudy (anonymous):

um so what other tests are there

OpenStudy (anonymous):

AST, p-series, root test, geometric series

OpenStudy (anonymous):

isn't my logic right lol? the series must be decreasing to be absolutely convergent, but it's increasing instead

OpenStudy (anonymous):

I'm not sure enough to say yes or no to that. See what happens when you do more conditionally convergent problems

OpenStudy (anonymous):

what about intergral test

OpenStudy (anonymous):

OH I think I got it

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

why integral test? we don't have any exponents lol

OpenStudy (anonymous):

nvm. My final guess is that its similar to the alternating harmonic series, which is absolutely convergent.

OpenStudy (anonymous):

but it's conditionally convergent lol

OpenStudy (anonymous):

\[\sqrt{k}\div x = 1/(k^1/2)\]

OpenStudy (anonymous):

K^(1/2)/K = 1/(k^(1/2))

OpenStudy (anonymous):

yeah idk this one is a feather

OpenStudy (anonymous):

>_< watch your language

OpenStudy (anonymous):

..serious?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

for series to converge they must be decreasing, but this one is increasing so that's why it's conditionally convergent ^_^

OpenStudy (anonymous):

right? .-.

OpenStudy (anonymous):

You can use the alternating series test.

OpenStudy (anonymous):

I did, but it didn't work >_<

OpenStudy (anonymous):

I ended up with 1 , which means no conclusion, the limit must be equal to 0

OpenStudy (anonymous):

You just need to show that sqrt(k)/(k+1) goes to zero as k goes to infinity (easy) and that sqrt(k)/(k+1) is monotonic decreasing. You can do this by taking the derivative and showing that the derivative is negative for all x greater than 1. \[f'(x)=\frac{1-x}{2\sqrt{x}(x+1)^2}=0 \rightarrow x=1\]partitions the interval. You take a test point x=4, say (anything in this interval) and show that \[f'(4)<0\]. It will be negative for all x in this interval (i.e x >= 1). By the Mean Value Theorem, you can choose any interval in [1, infinity), [a,b], say, and show that since f'(x)<0 on [1,infty), it is the case that\[f'(c)=\frac{f(b)-f(a)}{b-a}<0 \rightarrow f(b)<f(a)\]i.e. monotonic decreasing.

OpenStudy (anonymous):

but the limit is equal to = 1 loki

OpenStudy (anonymous):

\[\lim_{k \rightarrow \infty}\frac{k^{1/2}}{k+1}=\lim_{k \rightarrow \infty}\frac{1/k^{1/2}}{1+1/k}=0\]

OpenStudy (anonymous):

\[\lim_{k \rightarrow \infty} \frac{\sqrt(k)}{k+1} = \lim_{k \rightarrow \infty} \frac{|k|}{k} = 1\]

OpenStudy (anonymous):

No, it goes to zero...

OpenStudy (anonymous):

how come? .-.

OpenStudy (anonymous):

look above

OpenStudy (anonymous):

didn't I compute it right?

OpenStudy (anonymous):

I did lol

OpenStudy (anonymous):

I assume your numerator is\[\sqrt{k}\]

OpenStudy (anonymous):

there are alot of ways ^_^ you take the biggest power above, and the biggest power bellow then compute :)

OpenStudy (anonymous):

yep, which will be |k|........oh wait! LOL

OpenStudy (anonymous):

>_< silly silly mistake, k^1/2 - K = 1/k = 0 you're right, I'm sorry ^_^"

OpenStudy (anonymous):

so it's limit = 0 and it's decreasing using the ratio test/finding the derivative, so by AST it's convergent?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

but the book says it's conditionally convergent?

OpenStudy (anonymous):

that's the confusing part

OpenStudy (anonymous):

Well, the alternating series test is dependent upon the series having that (-)^k term in there. If it didn't, the sum wouldn't be killed off enough at every second step.

OpenStudy (anonymous):

Absolute convergence would have sqrt(k)/(k+1) as your series, not (-)^k etc...different.

OpenStudy (anonymous):

Actually, wait...

OpenStudy (anonymous):

hmm, the theorem says the following: if sigma (ak) is convergent, but sigma |ak| is divergent, then sigma|ak| is conditionally convergent. but using AST, both are convergent, which means that ak is absolutely convergent

OpenStudy (anonymous):

or wait! LOL you can use limit comparison test >_< where : \[ak = \frac{\sqrt(k)}{k+1}\] and\[bk = \frac{1}{\sqrt(k)}\]

OpenStudy (anonymous):

You can't use alternating series test on something that isn't alternating.

OpenStudy (anonymous):

|ak| isn't alternating.

OpenStudy (anonymous):

lol yep! so we can use limit comparison test ! ^_^

OpenStudy (anonymous):

then we must find limit of ak/bk

OpenStudy (anonymous):

right? :)

OpenStudy (anonymous):

I think you're being crazy. Use the alternating series test.

OpenStudy (anonymous):

Truuuuuuuuuuuusssssssssssssssttttttttttttttt mmmmmmmmmmmmmmmeeeeeeeeeeeeeeeeeeeee!!!!!!!!!!!!!!!! :)

OpenStudy (anonymous):

>_< lol, but AST gives absolute convergence, there's something wrong

OpenStudy (anonymous):

So are you now saying you want to test for convergence of |ak|?

OpenStudy (anonymous):

when I find the limit of ak/bk = 1 > 0, so it diverges hmmmm

OpenStudy (anonymous):

the question wants you to determine whether the series is absolutely convergent, conditionally convergent or divergent.

OpenStudy (anonymous):

wait limit comparison test won't work since it says if ak converges, then bk must converge and the vice versa, since they are a couple<- as pointed out by my professor lol

OpenStudy (anonymous):

WAIT! I GOT IT

OpenStudy (anonymous):

P -SERIES LOL

OpenStudy (anonymous):

Okay, if it's not abs. convergent, but convergent otherwise, it's conditionally convergent. If it's not convergent for either, it's divergent. We've shown it's convergent as is, now we show that it is divergent for |ak|...it will therefore be conditionally convergent.

OpenStudy (anonymous):

You can use the integral test for the absolute case.

OpenStudy (anonymous):

remember when I said that ak = swrt(k)/k+1 and bk = 1/sqrt(k)? bk is divergent as p-series, 1/2 < 1 so it's divergent, but ak is convergent, so it's conditionally convergent?

OpenStudy (anonymous):

why integral test? we use the integral test if we face exponentials and etc

OpenStudy (anonymous):

nah, it won't work, it's either both converge, or both diverge

OpenStudy (anonymous):

have mercy T_T

OpenStudy (anonymous):

hold on, can't we use root test? ^_^

OpenStudy (anonymous):

You can use limit comparison, with what you have. I got limit =1. It's greater than zero and finite, and since 1/sqrt(k) is divergent, |ak| is divergent too.

OpenStudy (anonymous):

but ak is convergent, so that's why it's conditonally convergent!

OpenStudy (anonymous):

LOKI, you're a genius LOL

OpenStudy (anonymous):

You're finished. 1) Limit comparison using 1/sqrt{k} to show |ak| divergent 2) Alternating series to show ak is convergent 3) Therefore, the series is conditionally convergent.

OpenStudy (anonymous):

o_o why use both ways?

OpenStudy (anonymous):

you were right just now >_<

OpenStudy (anonymous):

abt the limit comparison test lol

OpenStudy (anonymous):

You have to use both ways. Don't type anything until I've finished...

OpenStudy (anonymous):

okay, I'll behave ._.

OpenStudy (anonymous):

Absolute convergence is a stronger deal than conditional convergence. If a series is absolutely convergent, it is conditionally convergent. So abs. conv. is a *sufficient* condition to show convergence. If a series is not absolutely convergent, it doesn't mean it is necessary divergent. It may still be convergent. If it is, and does not have absolute convergence, you have *conditional convergence*. So people attack these problems first by looking for absolute convergence...why? Because the tests are easy and abs. convergence guarantees the other convergence. If it isn't abs. convergent, you STILL have to check for convergence.

OpenStudy (anonymous):

You can write now :D

OpenStudy (anonymous):

so, you're saying that absolute convergence = conditional convergence?

OpenStudy (anonymous):

= convergence?

OpenStudy (anonymous):

yes